8-1 Find the interest rates, or rates of return, on each of the following: a. You...

8-1

Find the interest rates, or rates of return, on each of the following:

a. You

borrow

$700 and promise to pay back $749 at the end of one year.

b. You

lend

$700 and receive a promise to be paid $749 at the end of one year.

c. You

borrow

$85,000 and promise to pay back $201,229 at the end of 10 years.

d. You

borrow

$9,000 and promise to make payments of $2,684.80 per year for five

years.
Answer #1

The rate is calculated using the RATE function as follows:

=RATE(nper,pmt,pv,fv)

A. =RATE(1,0,700,-749) = 7.00%

B. =RATE(1,0,-700,749) = 7.00%

C. =RATE(10,0,85000,-201229) = 9.00%

D. =RATE(5,-2684.8,9000) =15.00%